In a triangle a^2+b^2+c^2-ac-sqrt3*ab=0 then the triangle is

In a triangle
a^2+b^2+c^2-ac-sqrt3*ab=0
then the triangle is

Grade:9

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\a^2+b^2+c^2-ac-root3 *ab=0 \\$using quadratic formulae$ \\a={c+b\sqrt3 \pm \sqrt{-b^2-3c^2+2bc\sqrt3} \over 2} \\a={c+b\sqrt3 \pm \sqrt{-(b-c\sqrt3)^2} \over 2} \\=>b=c\sqrt3 \\=>a=2c \\ {a \over 1}={b \over \sqrt3}={c \over 1} \\=>sinA=1
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Leo Pixie
9 years ago
Great method